OpenStudy (anonymous):

Find all the factors of 110. Please show all of your work

OpenStudy (anonymous):

1,2,5,10,11,22,55

OpenStudy (anonymous):

as well as the negative

OpenStudy (anonymous):

-1 , -2 , -5, -10, -11, -22, -55

OpenStudy (anonymous):

The number of factors is based on the prime factorization. For instance, if x = p^n, where p is prime, then it has n+1 factors. If x = p^n * q^m, where p and q are prime, then it has (n + 1)(m + 1) factors We can continue as far as we need. 36 = 2^2*3^2, so it has (2 + 1)(2 + 1) = 3*3 = 9 factors. (To check:1, 2, 3, 4, 6, 9, 12 18, 36)

OpenStudy (anonymous):

thanx